OpenStudy (anonymous):

Simplify, showing your steps. Tell the property you used in each step. (3/2x)(2/3y) Simplify, showing your steps. Tell the property you used in each step. (3/2x)(2/3y) @Mathematics

OpenStudy (anonymous):

1/xy

OpenStudy (anonymous):

\[\frac{3}{2x} * \frac{2}{3y}\] =\[3*(2x)^{-1} * 2 * (3y)^{-1} \ \ \ \ \text{just rewriting stuff...} a/b = a * b^{-1}\] \[= 3 * 2^{-1}*x^{-1} * 2 * 3 ^{-1} * y^{-1}\ \ \ \ (ab)^{-1} = a^{-1}b^{-1}; \text{a and b are non-zero integers}\] \[=3*3^{-1} * 2 * 2^{-1} * x^{-1} * y^{-1} \ \ \ \ \text{commutative property of multiplication}\] \[=x^{-1} * y^{-1}\ \ \ \ \ \text{property of multiplicative inverse}\]then just rewrite that as \[\frac{1}{xy}\]

OpenStudy (anonymous):

you can go practice your skills on my next one? haha, I did it, just double checking my answers